The length of a helical spring is increased by 2 feet as a result of the suspension of a gravity of 8 pounds from its end
If the resistance of the medium in which the movement is taking is numerically equal to twice the instantaneous velocity, find the equation of motion if you know that the gravity seemed to move from the position of the equilibrium heading upward and at an initial velocity of 3 feet / s
Answer #1

c=216-s mg = 8 db 1.x = 28€ ag=32.17-ft 52 type of system, c²4 km = c2_45 malm + 2**= 24 c²4 km = 0.0216370 Cat is over dampe